Kindly solve this for me. I just want to know the value of y

y=100+0.3y+2000-2.1r
1
Expert's answer
2017-04-12T10:36:05-0400
If y = 100 + 0.3y + 2000 - 2.1r, then:
y - 0.3y = 2100 - 2.1r,
0.7y = 2100 - 2.1r,
y = 3000 - 3r.